and sales for the year total $51,850,000. Bad Debt Expense is estimated at 1/4 of 1% of sales. b. Determine the adjusted balances of Accounts Receivable, Allowance for Doubiful Accounts, and Bad Debt Expense Accounts Receivable Allowance for Doubtful Accounts Bad Debt Expense 110.000 1,205,000 X
Answer #1
a
Amount of adjusting entry 129625 = 51850000*1/4*1%
b
Accounts Receivable 2150000
Allowance for Doubtful accounts 119125 =129625-10500
Bad debt expense 129625
c
Net realizable value of Accounts Receivable 2030875 =2150000-119125